# Two algebraic proofs of the transcendence of $\mathrm{e}$ based on formal power series

### v1

@@ Abstract
We first remind the classical analytical proof of the transcendence of $\mathrm{e}$ due to Hilbert. Then, using formal power series, we give two algebraic semiformal proofs of the same result. The first proof is a specialization of a proof of the Lindemann– Weierstrass theorem given by Beukers, Bézivin and Roba [2]. The second proof uses improper integrals of formal power series and is due to this author. At the beginning, we explain what “semiformal” means and indicate yet another analytical proof of the transcendence of $\mathrm{e}$. Like the two algebraic proofs, it avoids uncountable sets. It is due to this author in [4].
@@
Hilbert’s proof [3] of the transcendence of the number $\mathrm{e}$ goes as follows. We suppose for the contradiction that there exists $n+1$, $n\in\mathbb{N}$ ($\{1,2,\dots\}$), integers $a_{0}$, $a_{1}$, $\dots$, $a_{n}$, not all of them 0, such that
@@
with $n$ as in $P$. We have the following expansions starting from the lowest powers of $x$:
$p_{r}(x)=(-1)^{n(r+1)}(n!)^{r+1}\cdot x^{r}+\bullet\cdot x^{r+1}+\dots$
@@
$p_{r}(x+j)=\bullet\cdot x^{r+1}+\bullet\cdot x^{r+2}+\dots$
for $j=1,2,\dots,n$ and some $\bullet\in\mathbb{Z}$ (the integers). We consider integrals
@@
Using additivity of the improper integral $\int_{0}^{+\infty}$ we express the product $P\cdot I_{r}$ as
$0$ $=$ $P\cdot I_{r}=A_{r}+B_{r}$
@@
where $c>1$ is a constant depending only on the coefficients $a_{0}$, $a_{1}$, $\dots$, $a_{n}$.
The heart of the proof is the determination of the second quantity
@@
We claim that $B_{r}$ is an integral multiple of $r!$ and that $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$. Since $c^{r}/r!\to 0$ as $r\to\infty$ for any constant $c>1$, the above equalities
@@
are contradictory and we are done. In order to prove the claim, we use the integral identity that
@@
for every $n\in\mathbb{N}_{0}$ ($\{0,1,\dots\}$). It is due to Euler and is easy to establish by induction on $n$ via integration by parts. By the linearity of the improper integral $\int_{0}^{+\infty}$, for any integers $b_{0}$, $b_{1}$, $\dots$, $b_{m}$ with $m\in\mathbb{N}_{0}$ we have more generally that
@@
for some $m_{r}\in\mathbb{Z}$. In the second line, we used the identity that $\mathrm{e}^{a+b}=\mathrm{e}^{a}\mathrm{e}^{b}$. In the third line, we changed the variable to y=xjy=x-j. In the fourth line, we used the above generalization of Euler’s identity and the expansions of $p_{r}(x+j)$, $j=0,1,\dots,n$, in lowest powers of $x$. It is clear that $B_{r}\in\mathbb{Z}$ and is divisible by $r!$. If $B_{r}=0$ then $(r+1)\,|\,a_{0}(n!)^{r+1}$. This divisibility is excluded when $r+1$ is coprime to the number $a_{0}\cdot n!$. Hence $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$. $\Box$
Interestingly, Baker [1, Theorem 1.2 on p. 4] gives a variant of the previous proof that uses only proper integrals $\int_{a}^{b}$.
Hilbert’s proof is elegant and short, but it is not completely satisfactory because it employs uncountable sets; such as, for example, the integrands in Euler’s identity. We explain how to fix it and how to prove the transcendence of the number $\mathrm{e}$ so that two conditions are satisfied.
@@
A set $x$ is HMC if for every index $n\in\mathbb{N}_{0}$ and every chain of sets
@@
the set $x_{n}$ is finite or countable. As for uses of mathematical logic, we quote a referee of [4]: “That basic real analysis is essentially a countable endeavor has been well-known for over 100 years. My secondhand understanding is that Weyl articulated this very well in Das Kontinuum, which nowadays we would say developed analysis in $\mathrm{ACA}_{0}$.” The reader can check [5] for $\mathrm{ACA}_{0}$. Our more demanding goal is to obtain a proof of the transcendence of $\mathrm{e}$ satisfying besides (A) also (B), that is, such that no HMC set and no mathematical logic is used.
There are (at least) three ways to achieve this goal. The first way is to replace real functions of the type $f\colon[a,b]\to\mathbb{R}$, where $a<b$ are real numbers and $\mathbb{R}$ is the set of real numbers, with the functions
@@
where $\mathbb{Q}$ is the (HMC) set of fractions. If real numbers are represented by Dedekind cuts, any such function $f$ is an HMC set. We realized the first way in [4]. Is there any difference in treatment of the two kinds of functions? The above mentioned referee thinks not: “Continuous functions on the reals are determined by their values on the rationals, so everything works out in the usual way. I don’t see how what the article is doing is so very different from developing analysis in (full) second-order arithmetic.” There is an obvious meta-argument refuting the referee’s first claim: if it were the case that “everything works out in the usual way,” then some enthusiasts would be teaching courses in analysis based on functions of the type $f\colon[a,b]\cap\mathbb{Q}\to\mathbb{R}$ for years. Since you have not heard about such courses, there must be some difficulties with this approach. Indeed, there are, but they can be overcome. The interested reader is referred to [4] for details. We include the referee’s second claim just to reiterate that our approach in [4] satisfies condition (B). In [4], we do not use technical notions of mathematical logic at all; therefore, our approach is very different from any countable treatment of real analysis offered by mathematical logic, such as, to give besides Das Kontinuum more modern references, [7] or Weaver’s forthcoming book [8].
We included this analytical digression to place the algebraic results of the present article in context. We present two algebraic proofs of the transcendence of $\mathrm{e}$ satisfying conditions (A) and (B). They are based on algebraic manipulations with formal power series. The first proof in Section 2 is a specialization of the proof in [2]. The second proof in Section 3 is due to this author. We presented it in early versions of [4], but we rework it here.
The word “directly” in condition (A) means that non-HMC sets may appear in the proof only by mention, as “harbors” of the main players. We may say, for example, that for every two real sequences $(a_{n}),(b_{n})\subset\mathbb{R}$ we have $\dots$; it is clear that $\mathbb{R}$ is used only indirectly and can be easily formulated away. We allow such uses of non-HMC sets for the sake of natural and easily understandable formulations.
In conclusion of this introduction, we review formal power series and explain the meaning of the adjective “semiformal”. If $R$ is an (integral) domain, we write $R[[x]]$ for the (usually non-HMC) set of maps $f\colon\mathbb{N}_{0}\to R$. They are traditionally written as infinite formal sums
$f=f(x)=\sum_{n\geq 0}a_{n}x^{n}$}\,,
where $a_{n}=f(n)$. We denote the coefficient $a_{n}$ also by $x^{n}]f$. The set $R[[x]]$ is endowed with the operations
${\textstyle\sum_{n\geq 0}a_{n}x^{n}+\sum_{n\geq 0}a_{n}x^{n}}$ $=$ ${\textstyle\sum_{n\geq 0}(a_{n}+b_{n})x^{n}\,\text{ and}}$
${\textstyle\sum_{n\geq 0}a_{n}x^{n}\cdot\sum_{n\geq 0}a_{n}x^{n}}$ $=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{k=0}^{n}a_{k}b_{n-k}\big)x^{n}\,,}$
@@
is a domain and $f$ is an $R[[x]]$-unit iff $x^{0}]f$ is an $R$-unit. An important unary operation on $R[[x]]$ is (formal) derivative,
@@
So for derivatives the characteristic of $R$ becomes relevant. Less often one uses the (formal) primitive
$\int\sum_{n\geq 0}a_{n}x^{n}=\sum_{n\geq 0}\frac{a_{n}}{n+1}x^{n+1}=\sum_{n\geq 1}\frac{a_{n-1}}{n}x^{n}\,$,}
but in this article, it is more important than the derivative. If $K$ is the normed field $K=\mathbb{R}$ or $K=\mathbb{C}$ (complex numbers), we denote by $K\{x\}$ the subdomain of $K[[x]]$ formed by convergent formal power series $f(x)=\sum_{n\geq 0}a_{n}x_{n}$ that satisfy
$\lim_{n\to\infty}a_{n}c^{n}=0$
for every $c\in K$. In other words, $f(x)$ has an infinite radius of convergence.
The binary and unary operations on $R[[x]]$ that we have listed so far, ++, $\cdot$, $(\dots)^{\prime}$, and $\int(\dots)$, are formal in the sense that, for example, $x_{n}](f\cdot g)$ is determined by finitely many coefficients in $f$ and $g$, and similarly for the other operations. In formal arguments involving elements $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $K[[x]]$ it is also forbidden to use values
$f(c)=\sum_{n\geq 0}a_{n}c^{n}=\lim_{n\to\infty}\sum_{k=0}^{n}a_{k}c^{k}$}
of $f(x)$ on elements $c\in K$. It is still an open problem to find a fully formal proof of the transcendence of $\mathrm{e}$ based on formal power series; or at least no such proof is known to this author. However, in the next two sections we present two such proofs if formality is relaxed to quasiformality.
An argument involving elements of $K[[x]]$ is quasiformal if it uses only formal and quasiformal operations, and for every $f(x)\in K[[x]]$ appearing in it only at most countably many values $f(c)$ for $c\in K$ are used. Quasiformal operations allow uses of limit transitions in $K$ and of at most countably many values $f(c)$.
Quasiformal operations on $\mathbb{R}[[x]]$ are introduced, for example, in Definitions 3.3 and 3.11.
@@ 2 Specializing the proof of Beukers, Bézivin and Robba
$b_{1}\mathrm{e}^{a_{1}}+b_{2}\mathrm{e}^{a_{2}}+\dots+b_{t}\mathrm{e}^{a_{t}}\neq 0$
whenever $b_{j}$ and $\alpha_{j}$ are $2t$, with $t\in\mathbb{N}$, algebraic numbers in $\mathbb{C}$ such that every $b_{j}\neq 0$ and the $\alpha_{j}$ are mutually distinct. See Baker [1, Theorem 1.4 on pp. 6–8] for an analytic proof. Using formal power series, Beukers, Bézivin and Robba gave in [2] another proof of the LW theorem. In this section, we specialize their proof to $b_{j},\alpha_{j}\in\mathbb{Z}$, which gives the transcendence of $\mathrm{e}$. In order to clearly describe logical structure of the proof, we completely reorganize the exposition in [2]. At the end of this section, we explain why this proof is semiformal (and not formal), and why it satisfies conditions (A) and (B).
We begin by recalling that a formal power series $f(x)$ in $\mathbb{C}[[x]]$ is rational if there exist two polynomials $p(x),q(x)\in\mathbb{C}[x]$ such that $q(0)=1$ and $q(x)f(x)=p(x)$; we may assume that $p(x)$ and $q(x)$ are coprime. If $f(x)$ and $g(x)$ in $\mathbb{C}[[x]]$ are rational, then so are $f(x)+g(x)$, $f(x)g(x)$, and $f^{\prime}(x)$. Any rational $f(x)\in\mathbb{C}[[x]]$ has a unique decomposition in partial fractions
@@
where $r(x)\in\mathbb{C}[x]$, $t\in\mathbb{N}_{0}$, $m_{j}\in\mathbb{N}$, $\beta_{j,i},\alpha_{j}\in\mathbb{C}$ with $\beta_{j,m_{j}}\neq 0$ and the $\alpha_{j}$ are nonzero and mutually distinct; for $t=0$, the double sum is interpreted as 0, and the numbers $m_{j}$ and $\alpha_{j}$ come from the factorization
@@
$\frac{\beta_{j,i}}{(1-\alpha_{j}x)^{i}}=\beta_{j,i}\sum_{n\geq 0}\binom{-i}{n}\alpha_{j}^{n}x^{n}\,.$}
@@
Proof. This follows from the definition of the decomposition of a rational formal power series in partial fractions. The product $p(x)A(x)$ possibly only decreases the orders of the poles of $A(x)$, or cancels them and does not create any new poles. The product $cx^{m}A^{\prime}(x)$ preserves all poles of $A(x)$, increases their orders by $1$, and does not create any new poles. The sum $p(x)A(x)+cx^{m}A^{\prime}(x)$ merges the two sets of poles. Therefore, the resulting set of poles with their orders is that of $cx^{m}A^{\prime}(x)$. $\Box$
We prove that $\mathrm{e}$ is transcendental by means of the following theorem on rationality of formal power series, which is extracted from the arguments in [2].
@@
We establish for $n,k\in\mathbb{N}_{0}$ with $n\geq tk$ two properties of the numbers $v_{n}(k)$.
@@
    We have bound $|v_{n}(k)|\leq cA^{n}C^{k}$ where $C=1+|a_{1}|+\dots+|a_{t}|$ and $c>0$ is a constant.
@@
The bound in item 1 follows by induction on $k$ and the assumed bound on $v_{n}=v_{n}(0)$. For $k=0$ the bound clearly holds. Let $n,k\in\mathbb{N}_{0}$ with $n\geq t(k+1)$. Then $n-t\geq tk$, and by the above recurrence and induction we have
@@
of the above recurrence, we show that for a fixed $k\in\mathbb{N}$ we have $v_{n}(k)=0$ for every large $n$. This will prove that the formal power series $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ is rational. So let $k_{0}\in\mathbb{N}$ be such that $k!>cA^{2tk}C^{k}$ for every $k\geq k_{0}$, and let
@@
From these three properties and from the modified recurrence, it follows by induction on the norm that $v_{n}(k)=0$ for every $\langle k,n\rangle\in M$. This concludes the proof because we see that every $k\geq k_{0}$ has the required property. $\Box$
@@
where $c>0$ is an absolute constant. Thus $v_{n}=O(A^{n})$ for $n\in\mathbb{N}_{0}$. Let $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ ($\in\mathbb{Z}[[x]]$). As we know from the previous proof, $v_{n}-nv_{n-1}=u_{n}$ for $n\in\mathbb{N}$ and $v_{0}=u_{0}$. In terms of formal power series we get
@@
We reiterate that the proof in [2] is more general as it establishes the full LW theorem. In this situation, one can only achieve that $u_{n},v_{n}\in\mathbb{Q}$, and denominators have to be taken care of. The previous proof works with formal power series, but it is only semiformal and not formal because the key player is the infinite sum
@@
which is the value $f(1)$ of the formal power series $f(x)=\sum_{n\geq 0}\frac{u_{n}}{n!}x^{n}$. In fact, this is the only non-formal feature of it.
It is clear that the previous proof satisfies condition (A), which states that all directly used sets are HMC. It works directly only with complex formal power series, which are maps from $\mathbb{N}_{0}$ to $\mathbb{C}$. As explained in Section 1, non-HMC sets like $\mathbb{C}[[x]]$ are not directly used in the proof. Condition (B), which states that no formalism of mathematical logic is used, is clearly satisfied as well. Although the article [2] of Beukers, Bézivin, and Robba is a pioneering one, as it is the first proof of the transcendence of the number $\mathrm{e}$ (in fact, of course, of the full LW theorem) that satisfies conditions (A) and (B), it is little known; in the database of Mathematical Reviews, it currently has only 11 citation (January 2026).

### v3

@@ Abstract
We remind the classical analytical proof of the transcendence of $\mathrm{e}$ due to Hilbert. Then, using formal power series, we give two algebraic semiformal proofs of this result. The first proof is a specialization of the proof of the Lindemann–Weierstrass theorem found by Beukers, Bézivin and Robba [2]. The second proof uses improper integrals of formal power series and is due to this author. We explain what “semiformal” means and indicate yet another analytical proof of the transcendence of $\mathrm{e}$. Like the two algebraic proofs, this proof avoids uncountable sets. It is due to this author in [7].
@@
The transcendence of the number $\mathrm{e}$ was first proven by Hermite [4]. Hilbert’s proof [5] goes as follows. We suppose for the contradiction that there exists $n+1$, $n\in\mathbb{N}$ ($\{1,2,\dots\}$), integers $a_{0}$, $a_{1}$, $\dots$, $a_{n}$, not all of them 0, such that
@@
with $n$ as in $P$. We have the expansions
$p_{r}(x)=(-1)^{n(r+1)}(n!)^{r+1}\cdot x^{r}+ax^{r+1}+\dots$
@@
$p_{r}(x+j)=bx^{r+1}+cx^{r+2}+\dots\,$,
where $j=1,2,\dots,n$ and $a,b,c,\dots$ are in $\mathbb{Z}$ (the integers). We consider improper integrals
@@
Using the additivity of improper integrals, we express the product $PI_{r}$ as
$0$ $=$ $PI_{r}=A_{r}+B_{r}$
@@
where $c\geq 1$ is a constant (depending only on the coefficients $a_{0}$, $a_{1}$, $\dots$, $a_{n}$).
We determine the second quantity
@@
We claim that $B_{r}$ is an integral multiple of $r!$ and that $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$. Since $c^{r}/r!\to 0$ as $r\to\infty$ for any constant $c\geq 1$, the above equalities
@@
are contradictory and we are done. We prove the claim by means of the integral identity that
@@
for every $n\in\mathbb{N}_{0}$ ($\{0,1,\dots\}$). The identity is due to Euler and is easy to establish by induction on $n$ via integration by parts. Using the linearity of improper integrals, we obtain for any integers $b_{0}$, $b_{1}$, $\dots$, $b_{m}$ with $m\in\mathbb{N}_{0}$ the more general identity that
@@
for some $m_{r}\in\mathbb{Z}$. In the second line, we used the identity that $\mathrm{e}^{a+b}=\mathrm{e}^{a}\mathrm{e}^{b}$. In the third line, we changed the variable to y=xjy=x-j. In the fourth line, we used the above generalization of Euler’s identity and the expansions of $p_{r}(x+j)$, $j=0,1,\dots,n$. It is clear that $B_{r}\in\mathbb{Z}$ and is divisible by $r!$. If $B_{r}=0$ then $(r+1)\,|\,a_{0}(n!)^{r+1}$. This divisibility is excluded when $r+1$ is coprime to the number $a_{0}\cdot n!$. Hence $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$. $\Box$
Baker [1, Theorem 1.2 on p. 4] gives a variant of the previous proof using only proper integrals $\int_{a}^{b}$. A similar proof is given in Conrad [3].
Hilbert’s proof is elegant and short, but it has the unsatisfactory feature that it employs uncountable sets, for example, the integrands in Euler’s identity. We explain how to improve upon it by proving the transcendence of $\mathrm{e}$ so that two conditions hold.
@@
We call proofs satisfying these two conditions AB proofs. A set $x$ is HMC if for every index $n\in\mathbb{N}_{0}$ and every chain of sets
@@
the set $x_{n}$ is at most countable (finite or countable). As for uses of mathematical logic in analysis, we quote a referee of [7]: “That basic real analysis is essentially a countable endeavor has been well-known for over 100 years. My secondhand understanding is that Weyl articulated this very well in Das Kontinuum, which nowadays we would say developed analysis in $\mathrm{ACA}_{0}$”. The reader can check [8] for $\mathrm{ACA}_{0}$. Our more ambitious and demanding goal is to obtain an AB proof of the transcendence of $\mathrm{e}$. In a related result, Kaibkhanov and Skopenkov [6] gave an AB proof of the transcendence of the Mahler number $\sum_{n\geq 0}2^{-2^{n}}$ based on binary expansions.
There are at least three ways to such proof. The first way is to replace real functions of the type $f\colon[a,b]\to\mathbb{R}$, where $a<b$ are real numbers and $\mathbb{R}$ is the set of real numbers, with the functions
@@
where $\mathbb{Q}$ is the set of fractions (if real numbers are Dedekind cuts, every such function $f$ is an HMC set), and to develop a sufficient fragment of real analysis for such functions. We realized this way in [7]. Is there any difference in the treatment of the two kinds of functions? The above mentioned referee thinks not: “Continuous functions on the reals are determined by their values on the rationals, so everything works out in the usual way. I don’t see how what the article is doing is so very different from developing analysis in (full) second-order arithmetic”. An obvious meta-argument refutes the first claim: if it were the case that “everything works out in the usual way”, then some enthusiasts would be teaching courses in analysis based on functions of the type $f\colon[a,b]\cap\mathbb{Q}\to\mathbb{R}$ for years. Since you have not heard about such courses, there must be some difficulties with this approach. Indeed, there are, but they can be overcome. The interested reader is referred to [7] for details. The second claim is refuted by the fact that in [7] we provide an AB proof. We do not use any technical notion of mathematical logic, and our approach in [7] is therefore very different from the countable treatments of real analysis offered by mathematical logic. For examples of the latter approaches, see Weaver [10, 11].
This analytical digression places the algebraic results of the present article in context. These results are two AB proofs of the transcendence of $\mathrm{e}$. Both are based on formal power series. The first proof in Section 2 is a specialization of the proof in [2]. The second proof in Section 3 is due to this author. We presented it already in early versions of [7], and we rework it here.
The word “directly” in condition (A) means that non-HMC sets may appear in the proof only as supersets of the main, directly used sets, which have to be HMC. We may say, for example, that for every two sequences $(a_{n}),(b_{n})\subset\mathbb{R}$ we have $\dots\;$. It is clear that the set $\mathbb{R}$ is mentioned only for convenience and can be easily formulated away. We allow such uses of non-HMC sets in AB proofs for the sake of natural and easily legible formulations.
In conclusion of the introduction, we review formal power series and explain the meaning of the adjective “semiformal”. If $R$ is an (integral) domain, we write $R[[x]]$ for the uncountable set of maps $f\colon\mathbb{N}_{0}\to R$. They are written as infinite formal sums
$f=f(x)=\sum_{n\geq 0}a_{n}x^{n}$}
with $a_{n}=f(n)$. We denote the coefficient $a_{n}$ also as $x^{n}]f$. The set $R[[x]]$ is endowed with the operations
${\textstyle f(x)+g(x)=\sum_{n\geq 0}a_{n}x^{n}+\sum_{n\geq 0}b_{n}x^{n}}$ $=$ ${\textstyle\sum_{n\geq 0}(a_{n}+b_{n})x^{n}\,\text{ and}}$
${\textstyle f(x)\cdot g(x)=\sum_{n\geq 0}a_{n}x^{n}\cdot\sum_{n\geq 0}b_{n}x^{n}}$ $=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{k=0}^{n}a_{k}b_{n-k}\big)x^{n}\,,}$
@@
is a domain and $f$ is an $R[[x]]$-unit iff $x^{0}]f$ is an $R$-unit. An important unary operation on $R[[x]]$ is the (formal) derivative,
@@
Less often one uses the (formal) primitive
$\int\sum_{n\geq 0}a_{n}x^{n}=\sum_{n\geq 0}\frac{a_{n}}{n+1}x^{n+1}=\sum_{n\geq 1}\frac{a_{n-1}}{n}x^{n}\,.$}
In this article, $\int$ is more important than the derivative. If $K$ is the normed field $K=\mathbb{R}$ or $K=\mathbb{C}$ (complex numbers), we denote by $K\{x\}$ the subdomain of $K[[x]]$ formed by the convergent formal power series $f(x)=\sum_{n\geq 0}a_{n}x_{n}$. They satisfy for every $c\in K$ that
$\lim_{n\to\infty}a_{n}c^{n}=0_{K}\,.$
In other words, $f(x)\in K\{x\}$ iff $f(x)$ has an infinite radius of convergence.
The binary and unary operations on $R[[x]]$ listed so far, ++, $\cdot$, $(\cdot)^{\prime}$, and $\int(\cdot)$, are formal in the sense that, for example, for every $n\in\mathbb{N}_{0}$ the coefficient
$x^{n}](f\cdot g)$
is determined by only finitely many coefficients in $f$ and $g$, and the same holds for the other operations. In formal proofs involving formal power series $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $K[[x]]$, only formal operations may be used, and it is forbidden to use the functional values
$f(c)=\sum_{n\geq 0}a_{n}c^{n}=\lim_{n\to\infty}\sum_{k=0}^{n}a_{k}c^{k},\ c\in K\,.$}
It is an open problem to give a completely formal AB proof of the transcendence of $\mathrm{e}$ based on formal power series. No such proof is known to this author.
However, we can present two such proofs in the next two sections if formality is relaxed to semiformality. Semiformal operations (in Section 3 we introduce several of them) are defined on $K\{x\}$, may use limit transitions, and may use at most countably many values $f(c)$ for $f\in K[[x]]$ and $c\in K$. Semiformal proofs concerning elements in $K[[x]]$ are proofs that satisfy the following condition.
Only formal and semiformal operations on $K[[x]]$ are used. For every formal power series involved in the proof, only at most countably many of its values are used.
In Sections 2 and 3 we give two AB &\& SF proofs of the transcendence of $\mathrm{e}$.
@@ 2 Specializing the proof of Beukers, Bézivin and Robba
$b_{1}\mathrm{e}^{\alpha_{1}}+b_{2}\mathrm{e}^{\alpha_{2}}+\dots+b_{t}\mathrm{e}^{\alpha_{t}}\neq 0$
whenever $b_{j}$ and $\alpha_{j}$ are $2t$, with $t\in\mathbb{N}$, algebraic numbers in $\mathbb{C}$ such that every $b_{j}\neq 0$ and the $\alpha_{j}$ are mutually distinct. See Baker [1, Theorem 1.4 on pp. 6–8] for an analytic proof. Using formal power series, Beukers, Bézivin, and Robba [2] obtained another proof of the LW theorem. In this section, we specialize their proof to $b_{j},\alpha_{j}\in\mathbb{Z}$, which gives the transcendence of $\mathrm{e}$. In order to clearly present the logical structure of their proof, we reorganize the exposition in [2].
We first recall that a formal power series $f(x)$ in $\mathbb{C}[[x]]$ is rational if there exist two polynomials $p(x),q(x)\in\mathbb{C}[x]$ such that $q(0)=1$ and $q(x)f(x)=p(x)$; we may assume that $p(x)$ and $q(x)$ are coprime. If $f(x)$ and $g(x)$ in $\mathbb{C}[[x]]$ are rational, then so are $f(x)+g(x)$, $f(x)g(x)$, and $f^{\prime}(x)$. Any rational formal power series $f(x)\in\mathbb{C}[[x]]$ has a unique decomposition into partial fractions
@@
where $r(x)\in\mathbb{C}[x]$, $t\in\mathbb{N}_{0}$, $m_{j}\in\mathbb{N}$, $\beta_{j,i},\alpha_{j}\in\mathbb{C}$ with $\beta_{j,m_{j}}\neq 0$, and the $\alpha_{j}$ are nonzero and mutually distinct. For $t=0$, the double sum is interpreted as 0, and the numbers $m_{j}$ and $\alpha_{j}$ come from the factorization
@@
$\frac{\beta_{j,i}}{(1-\alpha_{j}x)^{i}}=\sum_{n\geq 0}\binom{-i}{n}\beta_{j,i}\alpha_{j}^{n}\cdot x^{n}\,.$}
@@
Proof. The product $p(x)A(x)$ only possibly decreases the orders of the poles of $A(x)$ or cancels them and does not create any new poles. The product and derivative $cx^{m}A^{\prime}(x)$ preserve all poles of $A(x)$, increase their orders by $1$, and do not create any new poles. The sum $p(x)A(x)+cx^{m}A^{\prime}(x)$ merges the two sets of poles. It follows that the resulting set of poles with their orders is that of $cx^{m}A^{\prime}(x)$. $\Box$
We extracted the following theorem on rational formal power series from the arguments in [2].
@@
We establish for every $n,k\in\mathbb{N}_{0}$ with $n\geq tk$ two properties of the numbers $v_{n}(k)$.
@@
    We have the bound $|v_{n}(k)|\leq cA^{n}C^{k}$ where $C=1+|a_{1}|+\dots+|a_{t}|$ and $c\geq 0$ is a constant.
@@
The bound in item 1 follows by induction on $k$ from the assumed bound on $v_{n}=v_{n}(0)$. For $k=0$ the bound holds because it is the assumed bound on $v_{n}$. Let $n,k\in\mathbb{N}_{0}$ with $n\geq t(k+1)$. Then $n-t\geq tk$, and by the above recurrence and induction we have
@@
of the above recurrence, we show that there exists a $k\in\mathbb{N}$ such that $v_{n}(k)=0$ for every large $n$. This will prove that the formal power series $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ is rational. So let $k_{0}\in\mathbb{N}$ be such that $k!>cA^{2tk}C^{k}$ for every $k\geq k_{0}$, and let
@@
From these three properties, from the property of the set $N$ and from the modified recurrence, it follows by induction on the norm that $v_{n}(k)=0$ for every $\langle k,n\rangle\in M$. This concludes the proof because we see that every number $k\geq k_{0}$ has the required property. $\Box$
@@
where $c\geq 0$ is an absolute constant. Thus $v_{n}=O(A^{n})$ for $n\in\mathbb{N}_{0}$. Let $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ ($\in\mathbb{Z}[[x]]$). As we know from the previous proof, $v_{n}-nv_{n-1}=u_{n}$ for $n\in\mathbb{N}$ and $v_{0}=u_{0}$. In terms of formal power series we get
@@
We repeat that the proof in [2] is more general as it establishes the full LW theorem. In that proof, one can only achieve that $u_{n},v_{n}\in\mathbb{Q}$, and denominators have to be taken care of. The previous proof works with formal power series, but it is semiformal and not formal because the key player is the infinite sum
@@
which is the value $f(1)$ of the formal power series $f(x)=\sum_{n\geq 0}\frac{u_{n}}{n!}x^{n}$. In fact, this is the only non-formal component in an otherwise formal proof.
It is clear that the previous proof satisfies condition (A), which states that all directly used sets are HMC. It works directly only with complex formal power series, which are maps from $\mathbb{N}_{0}$ to $\mathbb{C}$. As explained in Section 1, non-HMC sets like $\mathbb{C}[[x]]$ are not really needed in the proof. Condition (B), which states that no formalism of mathematical logic is used, is satisfied as well. Although the article [2] of Beukers, Bézivin, and Robba is a pioneering one, as it is the first AB proof of the transcendence of the number $\mathrm{e}$ (in fact, of course, of the full LW theorem), it is little known; in the database of Mathematical Reviews, it currently has only 11 citation (January 2026).